Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

HELP plz! The probability tht an integrated circuit chip will have defective etching is 0.06, the probability tht it will have acrack defect is 0.03, and the probability tht it has both defects is 0.02 (a) what is the probability tht a newly manufactured chip will have either an etching or a crack defect? (b) what is the probability tht a newly manufactured chip will have neither defect?

OpenStudy (amistre64):

|dw:1428012049308:dw|

OpenStudy (amistre64):

P(AorB) = P(A) +P(B) -P(AandB) P(-A and -B) = 1 - P(AorB)

OpenStudy (anonymous):

Thanks amistre64 :)

OpenStudy (anonymous):

Amistre64 If the qts is tht fint the probability tht it will get one of the two defects dn???

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
22 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!